OpenStudy (lemanual):

Can someone please help me figure out what to do this is the directions... 1. Write an equation for the scenario given and use it to find the amount of flowers. 2. Identify is Max made an error in his work. If so, please explain what he did wrong and give the correct solution. 3. In the problem, they give you an equation. You need to use this equation to see how many adults purchased tickets if there were 5 children. 4. You can solve absolute value in order to find the max and min values.

OpenStudy (lemanual):

I'm suppose to be doing a grading rubric

OpenStudy (texaschic101):

what is the scenario given ??

OpenStudy (lemanual):

um idk :/ what does that mean

OpenStudy (texaschic101):

you can't make an equation without any information given

OpenStudy (lemanual):

Want me to show you the grading Rebric

OpenStudy (texaschic101):

I guess...this can't be the whole question

OpenStudy (lemanual):

Ok one sec

OpenStudy (lemanual):

Ok I found it

OpenStudy (lemanual):

Angelica's service club is selling flowers for a fundraiser. Each flower costs them $0.50 to purchase. They set the sale price of each flower at the purchase price plus 40%. Write and solve an equation to determine how many flowers they will need to sell to make a profit of $500. Use complete sentences to explain what each number and variable in your equation represents. Write your solution in the context of the original problem. Max was solving the equation below and he isn't sure if his answer is correct. Explain to Max how he can check his answer and then help him identify any errors he made. Provide the correct solution in your explanation. /Resource/22138797,A4C,0,C,42/Assets/problem.PNG Angelo took a group of people to the movies and spent $80 on tickets. Adult tickets cost $10, and child tickets cost $6. The equation to represent his cost is 10A + 6C = $80. If there were 5 children in the group, how many adult tickets did he purchase? Show your work and use a complete sentence to explain your solution in the context of the original problem. A quality control inspector is in charge of inspecting drill bits that are produced on an assembly line. The diameter of a drill bit can be modeled by the equation |x − 0.525| = 0.01. Show your work in solving the equation to determine the minimum and maximum diameters of the drill bits.

OpenStudy (lemanual):

I'm going to type the equation cause it wont let me do it.

OpenStudy (lemanual):

Here is the Equation

OpenStudy (texaschic101):

what equation did you post above ^^..where did that come from ?? sale price = purchase price + 40% sale price = 0.50 + .4(0.50) = 0.50 + 0.20 = 0.70 so the flowers are sold at 0.70 cents a piece which gives them a profit of .20 per flower they need to make a profit of $500 so your equation is : 0.20x = 500 -- (x represents # of flowers) x = 500/0.2 x = 2500 2500 flowers must be sold to make a profit of $500

OpenStudy (lemanual):

From the Instructions but i sent you the equation

OpenStudy (texaschic101):

let me check if the equation was solved correctly...give me a sec

OpenStudy (lemanual):

ok

OpenStudy (texaschic101):

mistake made... 3(3x + 6) = 4(7x - 1) 9x + 18 = 28x - 4 --- he distributed wrong here 9x - 9x + 18 = 28x - 9x - 4 18 = 19x - 4 18 + 4 = 19x 22 = 19x 22/19 = x

OpenStudy (texaschic101):

he put 28x - 1 and it should be 28x - 4

OpenStudy (lemanual):

Thanks so much :)

OpenStudy (lemanual):

Wait can you put that as a attach File so i can save that

OpenStudy (texaschic101):

I do not know how...I just worked it out on this post

OpenStudy (texaschic101):

I just do not see how that equation relates to the flower question

OpenStudy (lemanual):

Idk I just copy the E-mail my teacher sent me about what to do for the module one quiz

OpenStudy (lemanual):

Module two i meant

OpenStudy (texaschic101):

here is the problem about the tickets 10a + 6c = 80......a = adults, c = children If there were 5 children....so sub in 5 for c 10a + 6(5) = 80 10a + 30 = 80 10a = 80 - 30 10a = 50 a = 5 check.. 10a + 6c = 80 10(5) + 6(5) = 80 50 + 30 = 80 80 = 80 (correct) there were 5 adult tickets (a) and 5 childrens tickets (c) sold

OpenStudy (texaschic101):

last one.. | x - 0.525| = .01 x - 0.525 = .01 -(x - 0.525) = 0.01 x = .01 + 0.525 -x + 0.525 = 0.01 x = 0.535 -x = 0.01 - 0.525 -x = - 0.515 x = 0.515 minimum = 0.515 maximum = 0.535

OpenStudy (lemanual):

Thanks

OpenStudy (lemanual):

You're the best Ty so much. :D

OpenStudy (texaschic101):

YW :)